The Beetle Manufacturing Company produces the following three products:

Hammers Screwdriver Saws
Selling price per unit R40 R16 R50
Variable costs per unit R28 R12 R30
Contribution per unit R12 R 4 R20

Fixed costs are R76,000 per year.

50% of all sales in units are hammers, 30% are screwdrivers, and 20% are saws.


Required:

Calculate the following values:

A. break-even point in total units.

B. Number of hammers that will be sold at break-even.

C. Total sales in units to obtain a target income of R19,000.
Answer #1
Computation of Weighted Average Contribution ratio
Hammers Screwdriver Saws Total
Contribution Margin Per Unit (a) 12 4 20
Sales Mix (b) 50% 30% 20%
Weighted Average Contribution (axb) 6 1.2 4 11.2
a. BEP Unit= Fixed Cost / Weighted Average Contribution per Unit
=76000/11.20=6786 Unit
b. No. of Hammer to be sold
Total BEP Unit= 6786 Unit
Ratio of Hammer = 50%
No of Unit of Hammer ( 6786X50%)=3393 Unit
c. Tales Sales required for target Income
Revised Contribution = Fixed Cost + Target Income
=76000+19000=95000
Weighted Average Contribution = $11.20 per Unit
Total Sales in Units require= 95000/$11.20=8482 Unit